OpenStudy (anonymous):

cubic root of 9x^4 * cubic root of 7x^8

OpenStudy (nincompoop):

like this? \[\sqrt[3]{9x^4 } * \sqrt[3]{7x^8}\]

OpenStudy (nincompoop):

show your attempt at solving the problem

OpenStudy (anonymous):

\[\sqrt[3]{9x^4?}*\sqrt[3]{7x^8}\] \[\sqrt[3]{63x ^{12}}\] this as far as I could go, I have to simplify it

OpenStudy (bibby):

you can write roots as fractional exponents. for example:\[\huge \sqrt(x)=x^\frac{1}{2}\] \[\huge \sqrt{4}(x)=x^\frac{1}{4}\]

OpenStudy (bibby):

\[\huge \sqrt[4]{x}=x^\frac{1}{4}\] oopsies

OpenStudy (anonymous):

ok, so can you help solve this promblem

OpenStudy (bibby):

Apply the idea above to this:\[\huge \sqrt[3]{63x^{12}}\]

OpenStudy (anonymous):

so it would be \[63x ^{12}1/3\]

OpenStudy (bibby):

yeah. and what happens to exponents raised to other exponents?

OpenStudy (anonymous):

i dont know, we haven't learnt that

OpenStudy (bibby):

They're multiplied. \[\huge {x^{\frac{1}{5}}}^{25} = x^\frac{ 25 }{ 5 } = x^5\]

OpenStudy (anonymous):

so my problem would be \[x^4\]

OpenStudy (bibby):

you forgot the 63 I think

OpenStudy (anonymous):

yea i did so my answer would be 63x4?

OpenStudy (bibby):

I think so, yeah.

OpenStudy (anonymous):

ok thank you